You plant a 14 inch hemlock tree in your back yard that grows at a rate of 4 inches per year and an 8 inch blue spuce tree that grows at a rate of 6 inches per year. In how many years after you plant the trees will the two trees be the same height? And how tall will they be?
I have tried the following:
4x + 6y = 365
x + y = 2

Answer by scott8148(6628)

14 + 4y = 8 + 6y

6 = 2y ___ 3 = y

substituting ___ 14 + 4(3) = 26

it takes 3 yr and the trees are 26 in tall